# Model-Agnostic Policy Explanations with Large Language Models

### v1

@@ A.5.2 Instructions Given To Participants
For our IRB-approved participant studies, we create a website containing multiple choice problems for the participants to select. Prior to answering the problems, they are required to go through a slide deck where textual instructions and a tutorial is given. The tutorial is to help the participants understand our interface and ensure that they understand the mechanisms of the environment. The slide deck is provided as follows:
A.5.3 Slide Instructions
Figure 7: Slide 1
Figure 8: Slide 2
Figure 9: Slide 3
Figure 10: Slide 4
Figure 11: Slide 5
Figure 12: Slide 6
Figure 13: Slide 7
Figure 14: Slide 10
Figure 15: Slide 11
Figure 16: Slide 12
Figure 17: Slide 13
Figure 18: Slide 14
Figure 19: Slide 15

### v2

@@ A.5.2 Instructions Given To Participants
For our IRB-approved participant studies, we create a website containing multiple choice problems for the participants to select. Prior to answering the problems, they are required to go through a slide deck where textual instructions and a tutorial is given. The tutorial is to help the participants understand our interface and ensure that they understand the mechanisms of the environment.
The slide deck explains the roles and capabilities of the agents and basic knowledge about the USAR environment. Participants are then shown how natural language explanations are provided for each agent action and are instructed on how to assess these explanations for accuracy and helpfulness. The tutorial includes example scenarios and walks participants through two types of evaluation tasks: preference and next-action prediction.
A.6 Distillation Performance Analysis